OpenStudy (anonymous):

My answer is A :) What is the quotient: (6x3 + 7x2 - 30x - 25) ÷ (2x + 5) ? (3 points) Answer 3x2 + 4x - 5 3x2 - 4x - 5 3x2 - 4x + 5 3x2 + 4x + 5

OpenStudy (anonymous):

when you say 6x3 or 7x2 do you mean 6 to the third and 7 to the second?

OpenStudy (ja1):

IM pretty sure its b but let me make sure

OpenStudy (ja1):

In my opinion it's B but you would be taking a gamble with my opinion.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yes @RoseDryer

OpenStudy (anonymous):

Give me a minute to work it out and I'll let you know what I come up with

OpenStudy (anonymous):

okay thank you ;)

OpenStudy (anonymous):

gambling isn't always bad @JA1

OpenStudy (ja1):

true :D

OpenStudy (ja1):

maybe use mans best friend, google to help you

OpenStudy (anonymous):

B. just look at -25 5x? =-25 and \[7x^{2}\] at first all choice\[ \times 3x^{2}\] that mean 15x^2 and it need to be 7x^2

OpenStudy (anonymous):

I got C, I'm in trig btw if that helps your choice lol

OpenStudy (anonymous):

Here's my work: . . . . . ._ _ _ _ _ _ _ _ _ _ _ _ _ _ 2x + 5 | 6x³ + 7x² - 10x + 25 . . . . . - 6x³ - 15x² . . . . .. . . . .- 8x² - 10x . . . . .. . . . . .8x² + 20x . . . . .. . . . .. . . . .10x + 25 . . . . .. . . . .. . . .- 10x - 25 . . . . .. . . . .. . . . .. . . . .0

OpenStudy (anonymous):

hahaha thank you @RobZBHayes

OpenStudy (ja1):

yes i agree with rob

OpenStudy (anonymous):

I got: 3x^2 - (15/2)x + (29/4) with remainder -245/4

OpenStudy (anonymous):

Welcome!! Glad I could help

OpenStudy (anonymous):

you're the best. thanks everyone! xo

OpenStudy (anonymous):

oops umm... it's B. \[3x^{2}-4x-5\] so look \[3x^{2} \times (2x+5)=6x^{3}+15x^{2}\] \[-4x \times (2x+5)=-8x^{2}-20x\] \[-5 \times (2x+5)=-10x-25\] \[6x^{3}+15x^{2}-8x^{2}-20x-10x-25=6x^{3} + 7x^{2} - 30x - 25\] i think you can check this @InsanelyChaotic @JA1 @RobZBHayes i told you so

OpenStudy (anonymous):

nevermind if you still don't believe me :)

OpenStudy (anonymous):

thank you @DarthTony you're perf